# Cox reduction and confidence sets of models: a theoretical elucidation

### v2

@@ 1 Introduction
  • Arrange the $p$ variable indices at random either in a $k\times k$ square or a $k\times k\times k$ cube, where preferably $k\leq 15$. Extensions to four or more dimensions are possible. There is no loss of generality in assuming that $p$ is a perfect cube, as otherwise some positions are left unoccupied. A visualisation is in Supplementary Material Figure S1.
@@
A form of this procedure is presented in pseudo-code in Section S1 of the Supplementary Material.
@@
While inspiration came from [46], motivation for Cox reduction is from Bradford Hill’s [4] discussion of the circumstances under which an effect obtained in an observational study is relatively likely to have a causal interpretation. Such conditions include that the effect is reproduced in independent studies and behaves appropriately when the potential cause is applied, removed and then reinstated [13]. See [15, p.165–6] for further discussion. These aspects point to Cox reduction as at least a broadly appropriate approach for an initial phase of analysis, to be considered alongside some more established procedures such as marginal screening or penalised regression. We will show in Section 4 that the set of variables retained by Cox reduction is better suited than some other reduction strategies for assessment of models in the second phase. A further motivation for theoretical elucidation is that the procedure appears to be in use, as indicated by the download logs for the R package HCmodelSets [31]. Some published applications of Cox reduction include [3], [5], [33], [45].
@@
In a low-dimensional context, confidence sets of models have been emphasised repeatedly by Cox (see e.g., [12], [16], [17, Appendix A.2.5]). They have been considered from a different perspective in [30], which begins with a collection of candidate models and aims to identify those with the smallest expected loss at a given significance level. In [30], the relative explanatory power of pairs of models in the candidate set are compared, whereas [14] uses the comprehensive model as a reference set against which to gauge the adequacy of each submodel. If no submodel achieves comparable fit, the model confidence set of [14] is empty, whereas the set of [30] contains several models with equally poor fit. Bayesian approaches to model selection and the resulting credible sets of models are in the same vein, with obvious operational differences. For example, the posterior distributions in [37] and [29] assign non-zero weight to multiple models.
@@ 2 A simple explicit example
$Y=X\beta^{*}+\epsilon,\quad\quad\epsilon\sim N_{n}(\mathbf{0}_{n},\mathbb{I}_{n\times n})$ (1)
where $\beta^{*}=(1,1,0,\dots,0)^{T}\in\mathbb{R}^{p}$ with $p=25$ and $n=100$. Since $p$ is relatively small, there is no need for the reduction phase; confidence sets can be constructed directly. Rows of $X$ were generated independently from a normal distribution of zero mean and covariance
@@ 3 Notation
where the final equality follows from the proof of Lemma S5 in Section S5 of the supplementary material.
@@ 4.1 Framework
A model $\mathcal{S}_{m}$ is included in the model confidence set $\mathcal{M}$ if $\psi_{m}(Y,X)=1$. To reduce the computational burden, it is reasonable to construct the confidence set of models by only testing submodels of $\hat{\mathcal{S}}$ of size less than $s^{\#}$ say, chosen independently of $n$. The uncheckable assumption $s^{\#}\geq s$ is needed, otherwise the true model is necessarily excluded. Algorithm 1 in Section S1 of the supplementary material provides pseudo-code for the model assessment phase.
@@ 4.2 Behaviour under contiguous alternatives
Interpretation of the log-likelihood ratio test $\psi_{m}(Y,X)$ in general models is aided through appeal to Le Cam’s formulation of local asymptotic normality [34]. Under standard parametric regularity conditions, notably that the statistical model is differentiable in quadratic mean at $\gamma^{0}$, [44, Theorem 7.2] gives the following local asymptotic expansion of the log likelihood, which holds under model $\hat{\mathcal{S}}$ or any submodel thereof:
@@
In equation (3), $H_{0}$ is the Fisher information per observation, $Z_{n}=n^{-1/2}\nabla_{\gamma}\ell(\gamma^{0},0)$ converges in distribution to a normal random vector of mean zero and covariance matrix $H_{0}$, and $n^{1/2}h_{n}\rightarrow h$. Sufficient conditions for differentiability in quadratic mean are given by [44, Proposition 7.6]. These are satisfied, for example, by most exponential family regression models. It follows from (3) that
@@
Evans [23] generalised this result to allow two models, whose components are both in a $O(n^{-1/2})$ neighbourhood of the true distribution. This is a restriction to so-called contiguous alternatives, which in the context of equation (2) would imply that the model indexed by $\mathcal{S}_{m}$, if false, only omits variables whose associated signal strengths are dominated by the estimation error of the maximum likelihood estimator. This characterisation belongs to formal theory but indicates the directions from $\gamma^{0}$ in which the likelihood ratio test is expected to have relatively high or low power. As in [23], introduce $h_{n}$ and $\widetilde{h}_{n}$ such that $h_{n}\rightarrow 0$, $\widetilde{h}_{n}\rightarrow 0$ and $n^{1/2}(h_{n}-\widetilde{h}_{n})\rightarrow k$. Provided that the true distribution belongs to a model whose density with respect to an appropriate dominating measure is doubly differentiable in quadratic mean, in the sense of [23, Definition 2.8], then
@@ 4.3 Extensions to other alternatives
[…]
@@
where with slight notational inaccuracy, $\eta$ lies on the line joining $\eta^{0}+h$ and $\eta^{0}$, and may differ in each entry of the matrix. In certain generalised linear models, this term is of order $\|h\|_{3}^{3}$ which motivates assumptions (b) and (c). For further details, see the proof of Lemma S10 in the supplementary material.
Section S6 of the supplementary material gives conditions under which the assumptions in Proposition 1 are satisfied by canonical generalised linear models. These conditions are met by the linear and logistic regression models when $\mathcal{S}_{m}$ satisfies $\|(I-P_{m})\widetilde{X}\gamma^{0}\|_{2}=o(n^{-1/2})$ and so those models that only exclude a small portion of the signal $\widetilde{X}\gamma^{0}$ will be included in the model confidence set. It is not necessary for $\widetilde{X}$ to be full rank; see the proof of Proposition 1 for details. The expected size of the model confidence set is at least $M(1-\vartheta)$ where $M$ denotes the number of models $\mathcal{S}_{m}$ satisfying $\|(I-P_{m})\widetilde{X}\gamma^{0}\|=o(n^{-1/2})$. If there are many sets of covariates that are highly correlated in sample with signal variables, then the model confidence set will contain a large number of models on average.
@@ 5.1 Penalised regression
Penalised regression may also fail to select all signal variables even when they are highly correlated with the response variable. This occurrence is explained in [22] for the linear model with LASSO solution $\hat{\theta}^{(\lambda)}$ obtained using the LARS algorithm. Briefly, given a current predicted response, the algorithm sequentially adds the covariate that is most highly correlated with the residuals to update the predicted response. The sequence of solutions obtained at each step of the algorithm corresponds to the solutions of the LASSO problem for decreasing $\lambda$ [22, Theorem 1]. When two covariates are highly but not perfectly correlated in sample, including one of them in a LARS step reduces the correlation between the other and the residual, thus making it difficult for the second variable to enter the model. This can lead the LASSO to select unimportant noise variables over signal variables. The simulation results summarised at the start of Section 8 provide examples of this, showing that when signal strengths are weak and correlations among covariates are high, the comprehensive model determined by an undertuned LASSO often omits signal variables.
@@ 5.2 Marginal screening
where $\mathit{E},\mathit{F}\subset\{1,\dots,p\}$ with $|\mathit{E}|=1$. The result (6) was first noted in [10] for sets $\mathit{F}$ of size one and has been used subsequently in [11, 20]. A derivation for $|\mathit{F}|>1$ in the linear model is given in Section S4 of the supplementary material (proof of Lemma S1). An asymptotic analogue of equation (6) for general regression models was given by [18]. It follows from (6) that $\hat{\theta}_{Y:\mathit{E}}=\hat{\theta}_{Y:\mathit{E}.\mathit{F}}$ if and only if either $\hat{\theta}_{\mathit{F}:\mathit{E}}=0$ or $\hat{\theta}_{Y:\mathit{F}.\mathit{E}}=0$ or the two vectors are orthogonal. In particular, if this condition is violated when $\mathit{E}$ indexes a signal variable and $\mathit{F}=\mathcal{S}\backslash\mathit{E}$, it is possible for the marginal effect to be inflated or diminished. This improves or curtails the ability of marginal screening to detect $\mathit{E}$. If there is total cancellation, i.e. $\hat{\theta}_{Y:\mathit{E}.\mathit{F}}\approx-(\hat{\theta}_{\mathit{F}:\mathit{E}})^{T}\hat{\theta}_{Y:\mathit{F}.\mathit{E}}$, marginal screening would be unable to detect the signal variable indexed by $\mathit{E}$.
@@ 6.1 Framework
  • In the first round of Cox reduction, variable indices are randomly arranged in a $k\times k\times k$ cube. Variables are retained for further analysis if they are among the two most significant in at least two out of the three regressions in which they appear. See Algorithm 2 in Section S1 of the supplementary material.
  • In the second round of Cox reduction, the indices corresponding to variables retained after the first round are randomly arranged in a $k\times k$ square. The comprehensive model indexes those variables that are significant at level $\alpha$ in at least one regression. See Algorithm 3 in Section S1 of the supplementary material.
@@ 6.3 Preliminary insights
where $D_{\mathit{K}}$ is a diagonal matrix with entries given by the diagonal entries of $(X_{\mathit{K}}^{T}X_{\mathit{K}})^{-1}$ and $\hat{\theta}_{Y:\mathit{K}}$ was defined in Section 5.2. Lemma S2 shows that the entry of $T_{Y:\mathit{K}}$ corresponding to variable index $e\in\mathit{K}$ is
@@ 6.4 First reduction
Then the variable indexed by $e$ is retained through the first reduction if at least two of the three events $\mathcal{E}_{e:\mathit{K}_{1}}$, $\mathcal{E}_{e:\mathit{K}_{2}}$ and $\mathcal{E}_{e:\mathit{K}_{3}}$ occur. See Algorithm 2 in Section S1 of the Supplementary material.
@@ 6.5 Second reduction
Suppose that the index $e$, retained through the first reduction, is randomly arranged in a $k\times k$ square, where the dimension $k$ may be different from that of the first reduction. Let $\mathit{K}_{1}$ and $\mathit{K}_{2}$ be the variables that share a row or column with $e$ in the square and redefine $\mathcal{E}_{e:\mathit{K}}$ to be the event that the entry of $T_{Y:\mathit{K}}$ corresponding to $e\in\mathit{K}$ is significant at level $\alpha$. Then, the second round of reduction retains the variable indexed by $e$ on the event $\mathcal{E}_{e:\mathit{K}_{1}}\cup\mathcal{E}_{e:\mathit{K}_{2}}$. See Algorithm 3 in Section S1 of the Supplementary material.
@@
The second reduction requires any apparent explanatory power to persist when the variable in question is accompanied by relatively strong companion variables. Its intended purpose was to differentiate variables in $\mathcal{S}$ from those in $\mathcal{S}^{c}\cap\mathcal{A}$, when $\mathcal{S}\subseteq\mathcal{A}$. Consider a regression of $Y$ on $X_{\mathit{A}}$ where $\mathit{A}\subseteq\mathcal{A}$ contains signal and noise indices. Lemma S3 shows that the entry of $\sigma T_{Y:A}$ corresponding to $a\in\mathit{A}$ is given by
@@ 8 Numerical performance
Section S2 of the supplementary material shows the numerical performance of Cox reduction and the proposed construction of the model confidence set. Simulated data were generated to compare marginal screening and LASSO penalised regression to Cox reduction in the setting laid out in Section 6. Cox reduction and marginal screening both performed well, retaining all signal variables in the comprehensive model and yielding a model confidence set with high simulated coverage probability. In contrast, the undertuned LASSO failed to include all signal variables in the comprehensive model when correlations among variables were high and signal strengths were low. The size of the model confidence set constructed from marginal screening contained more models on average than that obtained from Cox reduction, as expected based on the discussion in Section 5. The recommended improvements to Cox reduction from Section 7 performed favourably.
@@
In each of 500 Monte Carlo replications an artificial response was generated anew and a confidence set of models was constructed using the most competitive of the reduction procedures from the analysis in Section S2 of the supplementary material, namely marginal screening (MS), Cox reduction with rerandomisation (CR-R) and Cox reduction with rerandomisation and sample splitting (CR-RSS). A set of indices $\mathcal{I}\subset\{1,\ldots,n\}$ of size $|\mathcal{I}|=29$ was drawn at random and fixed across Monte Carlo replications. This was used for model assessment for all procedures. The observations indexed by $\mathcal{I}^{c}$ were used for reduction for MS and CR-R, while CR-RSS used $\mathcal{I}$ for its first reduction and $\mathcal{I}^{c}$ for its second. For each approach, we recorded the proportion of times each signal variable was included in $\hat{\mathcal{S}}$, the proportion of times $\hat{\mathcal{S}}$ included the full set of signal variables, the proportion of times the set of signal variables was included in the model confidence set $\mathcal{M}$ and the size of $\mathcal{M}$.
@@
A favourable situation for both marginal screening and Cox reduction is when the estimated absolute effect appears larger than it is due to strong dependence between signal variables of an appropriate sign. This situation was explored and the results were similar to those reported in the analysis in Section S2 and are therefore omitted. The aim of the following three examples is to highlight limitations of the methodology in the presence of certain covariate dependencies.

### v3

@@ 1 Introduction
  • Arrange the $p$ variable indices at random either in a $k\times k$ square or a $k\times k\times k$ cube, where preferably $k\leq 15$. Extensions to four or more dimensions are possible. There is no loss of generality in assuming that $p$ is a perfect cube, as otherwise some positions are left unoccupied. A visualisation is in Figure S1 of the supplementary material [36].
@@
A form of this procedure is presented in pseudo-code in Section S1 of the supplementary material [36].
@@
While inspiration came from [47], motivation for Cox reduction is from Bradford Hill’s [30] discussion of the circumstances under which an effect obtained in an observational study is relatively likely to have a causal interpretation. Such conditions include that the effect is reproduced in independent studies and behaves appropriately when the potential cause is applied, removed and then reinstated [12]. See [14, p.165–6] for further discussion. These aspects point to Cox reduction as at least a broadly appropriate approach for an initial phase of analysis, to be considered alongside some more established procedures such as marginal screening or penalised regression. We will show in Section 4 that the set of variables retained by Cox reduction is better suited than some other reduction strategies for assessment of models in the second phase. A further motivation for theoretical elucidation is that the procedure appears to be in use, as indicated by the download logs for the R package HCmodelSets [31]. Some published applications of Cox reduction include [3], [4], [33], [39].
@@
In a low-dimensional context, confidence sets of models have been emphasised repeatedly by Cox (see e.g., [11], [15], [16, Appendix A.2.5]). They have been considered from a different perspective in [29], which begins with a collection of candidate models and aims to identify those with the smallest expected loss at a given significance level. In [29], the relative explanatory power of pairs of models in the candidate set are compared, whereas [13] uses the comprehensive model as a reference set against which to gauge the adequacy of each submodel. If no submodel achieves comparable fit, the model confidence set of [13] is empty, whereas the set of [29] contains several models with equally poor fit. Bayesian approaches to model selection and the resulting credible sets of models are in the same vein, with obvious operational differences. For example, the posterior distributions in [38] and [28] assign non-zero weight to multiple models.
@@ 2 A simple explicit example
$Y=X\theta^{0}+\epsilon,\quad\quad\epsilon\sim N_{n}(\mathbf{0}_{n},\mathbb{I}_{n\times n})$ (1)
where $\theta^{0}=(1,1,0,\dots,0)^{T}\in\mathbb{R}^{p}$ with $p=25$ and $n=100$. Since $p$ is relatively small, there is no need for the reduction phase; confidence sets can be constructed directly. Rows of $X$ were generated independently from a normal distribution of zero mean and covariance
@@ 3 Notation
where the final equality follows from the proof of Lemma S5 in [36].
@@ 4.1 Framework
A model $\mathcal{S}_{m}$ is included in the model confidence set $\mathcal{M}$ if $\psi_{m}(Y,X)=1$. To reduce the computational burden, it is reasonable to construct the confidence set of models by only testing submodels of $\hat{\mathcal{S}}$ of size less than $s^{\#}$ say, chosen independently of $n$. The uncheckable assumption $s^{\#}\geq s$ is needed, otherwise the true model is necessarily excluded. Algorithm 1 in Section S1 of the supplementary material of [36] provides pseudo-code for the model assessment phase.
@@ 4.2 Behaviour under contiguous alternatives
Interpretation of the log-likelihood ratio test $\psi_{m}(Y,X)$ in general models is aided through appeal to Le Cam’s formulation of local asymptotic normality [34]. Under standard parametric regularity conditions, notably that the statistical model is differentiable in quadratic mean at $\gamma^{0}$, [46, Theorem 7.2] gives the following local asymptotic expansion of the log likelihood, which holds under model $\hat{\mathcal{S}}$ or any submodel thereof:
@@
In equation (3), $H_{0}$ is the Fisher information per observation, $Z_{n}=n^{-1/2}\nabla_{\gamma}\ell(\gamma^{0},0)$ converges in distribution to a normal random vector of mean zero and covariance matrix $H_{0}$, and $n^{1/2}h_{n}\rightarrow h$. Sufficient conditions for differentiability in quadratic mean are given by [46, Proposition 7.6]. These are satisfied, for example, by most exponential family regression models. It follows from (3) that
@@
Evans [22] generalised this result to allow two models, whose components are both in a $O(n^{-1/2})$ neighbourhood of the true distribution. This is a restriction to so-called contiguous alternatives, which in the context of equation (2) would imply that the model indexed by $\mathcal{S}_{m}$, if false, only omits variables whose associated signal strengths are dominated by the estimation error of the maximum likelihood estimator. This characterisation belongs to formal theory but indicates the directions from $\gamma^{0}$ in which the likelihood ratio test is expected to have relatively high or low power. As in [22], introduce $h_{n}$ and $\widetilde{h}_{n}$ such that $h_{n}\rightarrow 0$, $\widetilde{h}_{n}\rightarrow 0$ and $n^{1/2}(h_{n}-\widetilde{h}_{n})\rightarrow k$. Provided that the true distribution belongs to a model whose density with respect to an appropriate dominating measure is doubly differentiable in quadratic mean, in the sense of [22, Definition 2.8], then
@@ 4.3 Extensions to other alternatives
$\hat{\eta}_{m}$ $=$ $\text{argmax}_{\eta\in\text{Col-Sp}(\widetilde{X}_{m})}\bar{\ell}(\eta)$
$\hat{\eta}$ $=$ $\text{argmax}_{\eta\in\text{Col-Sp}(\widetilde{X})}\bar{\ell}(\eta).$
@@
where with slight notational inaccuracy, $\eta$ lies on the line joining $\eta^{0}+h$ and $\eta^{0}$, and may differ in each entry of the matrix. In certain generalised linear models, this term is of order $\|h\|_{3}^{3}$ which motivates assumptions (b) and (c). For further details, see the proof of Lemma S10 in the supplementary material of [36].
Section S6 of the supplementary material of [36] gives conditions under which the assumptions in Proposition 1 are satisfied by canonical generalised linear models. These conditions are met by the linear and logistic regression models when $\mathcal{S}_{m}$ satisfies $\|(I-P_{m})\widetilde{X}\gamma^{0}\|_{2}=o(n^{-1/2})$ and so those models that only exclude a small portion of the signal $\widetilde{X}\gamma^{0}$ will be included in the model confidence set. It is not necessary for $\widetilde{X}$ to be full rank; see the proof of Proposition 1 for details. The expected size of the model confidence set is at least $M(1-\vartheta)$ where $M$ denotes the number of models $\mathcal{S}_{m}$ satisfying $\|(I-P_{m})\widetilde{X}\gamma^{0}\|=o(n^{-1/2})$. If there are many sets of covariates that are highly correlated in sample with signal variables, then the model confidence set will contain a large number of models on average.
@@ 5.1 Penalised regression
Penalised regression may also fail to select all signal variables even when they are highly correlated with the response variable. This occurrence is explained in [21] for the linear model with LASSO solution $\hat{\theta}^{(\lambda)}$ obtained using the LARS algorithm. Briefly, given a current predicted response, the algorithm sequentially adds the covariate that is most highly correlated with the residuals to update the predicted response. The sequence of solutions obtained at each step of the algorithm corresponds to the solutions of the LASSO problem for decreasing $\lambda$ [21, Theorem 1]. When two covariates are highly but not perfectly correlated in sample, including one of them in a LARS step reduces the correlation between the other and the residual, thus making it difficult for the second variable to enter the model. This can lead the LASSO to select unimportant noise variables over signal variables. The simulation results summarised at the start of Section 8 provide examples of this, showing that when signal strengths are weak and correlations among covariates are high, the comprehensive model determined by an undertuned LASSO often omits signal variables.
@@ 5.2 Marginal screening
where $\mathit{E},\mathit{F}\subset\{1,\dots,p\}$ with $|\mathit{E}|=1$. The result (6) was first noted in [9] for sets $\mathit{F}$ of size one and has been used subsequently in [10, 19]. A derivation for $|\mathit{F}|>1$ in the linear model is given in Section S4 of the supplementary material of [36] (proof of Lemma S1). An asymptotic analogue of equation (6) for general regression models was given by [17]. It follows from (6) that $\hat{\theta}_{Y:\mathit{E}}=\hat{\theta}_{Y:\mathit{E}.\mathit{F}}$ if and only if either $\hat{\theta}_{\mathit{F}:\mathit{E}}=0$ or $\hat{\theta}_{Y:\mathit{F}.\mathit{E}}=0$ or the two vectors are orthogonal. In particular, if this condition is violated when $\mathit{E}$ indexes a signal variable and $\mathit{F}=\mathcal{S}\backslash\mathit{E}$, it is possible for the marginal effect to be inflated or diminished. This improves or curtails the ability of marginal screening to detect $\mathit{E}$. If there is total cancellation, i.e. $\hat{\theta}_{Y:\mathit{E}.\mathit{F}}\approx-(\hat{\theta}_{\mathit{F}:\mathit{E}})^{T}\hat{\theta}_{Y:\mathit{F}.\mathit{E}}$, marginal screening would be unable to detect the signal variable indexed by $\mathit{E}$.
@@ 6.1 Framework
  • In the first round of Cox reduction, variable indices are randomly arranged in a $k\times k\times k$ cube. Variables are retained for further analysis if they are among the two most significant in at least two out of the three regressions in which they appear. See Algorithm 2 in Section S1 of the supplementary material [36].
  • In the second round of Cox reduction, the indices corresponding to variables retained after the first round are randomly arranged in a $k\times k$ square. The comprehensive model indexes those variables that are significant at level $\alpha$ in at least one regression. See Algorithm 3 in Section S1 of the supplementary material [36].
@@ 6.3 Preliminary insights
where $D_{\mathit{K}}$ is a diagonal matrix with entries given by the diagonal entries of $(X_{\mathit{K}}^{T}X_{\mathit{K}})^{-1}$ and $\hat{\theta}_{Y:\mathit{K}}$ was defined in Section 5.2. Lemma S2 in [36] shows that the entry of $T_{Y:\mathit{K}}$ corresponding to variable index $e\in\mathit{K}$ is
@@ 6.4 First reduction
Then the variable indexed by $e$ is retained through the first reduction if at least two of the three events $\mathcal{E}_{e:\mathit{K}_{1}}$, $\mathcal{E}_{e:\mathit{K}_{2}}$ and $\mathcal{E}_{e:\mathit{K}_{3}}$ occur. See Algorithm 2 in Section S1 of the supplementary material [36].
@@ 6.5 Second reduction
Suppose that the index $e$, retained through the first reduction, is randomly arranged in a $k\times k$ square, where the dimension $k$ may be different from that of the first reduction. Let $\mathit{K}_{1}$ and $\mathit{K}_{2}$ be the variables that share a row or column with $e$ in the square and redefine $\mathcal{E}_{e:\mathit{K}}$ to be the event that the entry of $T_{Y:\mathit{K}}$ corresponding to $e\in\mathit{K}$ is significant at level $\alpha$. Then, the second round of reduction retains the variable indexed by $e$ on the event $\mathcal{E}_{e:\mathit{K}_{1}}\cup\mathcal{E}_{e:\mathit{K}_{2}}$. See Algorithm 3 in Section S1 of the supplementary material [36].
@@
The second reduction requires any apparent explanatory power to persist when the variable in question is accompanied by relatively strong companion variables. Its intended purpose was to differentiate variables in $\mathcal{S}$ from those in $\mathcal{S}^{c}\cap\mathcal{A}$, when $\mathcal{S}\subseteq\mathcal{A}$. Consider a regression of $Y$ on $X_{\mathit{A}}$ where $\mathit{A}\subseteq\mathcal{A}$ contains signal and noise indices. Lemma S3 in [36] shows that the entry of $\sigma T_{Y:A}$ corresponding to $a\in\mathit{A}$ is given by
@@ 8 Numerical performance
Section S2 of the supplementary material [36] shows the numerical performance of Cox reduction and the proposed construction of the model confidence set. Simulated data were generated to compare marginal screening and LASSO penalised regression to Cox reduction in the setting laid out in Section 6. Cox reduction and marginal screening both performed well, retaining all signal variables in the comprehensive model and yielding a model confidence set with high simulated coverage probability. In contrast, the undertuned LASSO failed to include all signal variables in the comprehensive model when correlations among variables were high and signal strengths were low. The size of the model confidence set constructed from marginal screening contained more models on average than that obtained from Cox reduction, as expected based on the discussion in Section 5. The recommended improvements to Cox reduction from Section 7 performed favourably.
@@
In each of 500 Monte Carlo replications an artificial response was generated anew and a confidence set of models was constructed using the most competitive of the reduction procedures from the analysis in Section S2 of the supplementary material [36], namely marginal screening (MS), Cox reduction with rerandomisation (CR-R) and Cox reduction with rerandomisation and sample splitting (CR-RSS). A set of indices $\mathcal{I}\subset\{1,\ldots,n\}$ of size $|\mathcal{I}|=29$ was drawn at random and fixed across Monte Carlo replications. This was used for model assessment for all procedures. The observations indexed by $\mathcal{I}^{c}$ were used for reduction for MS and CR-R, while CR-RSS used $\mathcal{I}$ for its first reduction and $\mathcal{I}^{c}$ for its second. For each approach, we recorded the proportion of times each signal variable was included in $\hat{\mathcal{S}}$, the proportion of times $\hat{\mathcal{S}}$ included the full set of signal variables, the proportion of times the set of signal variables was included in the model confidence set $\mathcal{M}$ and the size of $\mathcal{M}$.
@@
A favourable situation for both marginal screening and Cox reduction is when the estimated absolute effect appears larger than it is due to strong dependence between signal variables of an appropriate sign. This situation was explored and the results were similar to those reported in the analysis in Section S2 in [36] and are therefore omitted. The aim of the following three examples is to highlight limitations of the methodology in the presence of certain covariate dependencies.